High Desert Potterworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costingWhat would be the total manufacturing cost recorded for Job 205? (Round Predetermined overhead rate to 2 decimal places. RoIf the job contained 28 units, what would be the unit product cost? (Round Predetermined overhead rate to 2 decimal places.
Answer #1
Predetermined overhead rate:
Molding 4.90 =2.8+(170100/81000)
Painting 13.70 =4.8+(492170/55300)
Overhead applied:
Molding 1813 =370*4.90
Painting 1740 =127*13.70
1
Department
Molding Painting Total
Direct materials 948 1280 2228
Direct labor 740 990 1730
Manufacturing Overhead 1813 1740 3553
Total cost 3501 4010 7511
2
Total cost 7511
Divide by units 28
Unit product cost 268.25 per unit